In how many ways can a committee of three men and four women be formed from a group of 8 men and 8 women?
a. 564480 b. 3920 c. 560 d. 70 e. None of the abo

Answer by ewatrrr(24785)

 
Hi
In how many ways can a committee of three men and four women be formed from a
group of 8 men and 8 women?
8C3*8C4 = 56*70 = 3920 ways